OpenStudy (anonymous):

The equation below shows the perimeter f(x) of several rectangular pieces of cardboard having width x inches: f(x) = 52 + 2x The perimeter of each piece is less than 100 inches. Which of the following best describes the domain of f(x)? All real numbers greater than 0 and less than 100 All real numbers greater than 0 and less than 24 All integers greater than 0 and less than 100 All integers greater than 0 and less than 24

OpenStudy (sleepyjess):

Is there anything else that you could eliminate just by looking and not even solving anything?

OpenStudy (anonymous):

Well integers are negative and positive numbers including zero. But integers don't include fractions or decimals so I don't think it's A or B

OpenStudy (sleepyjess):

So you think it is integers?

OpenStudy (anonymous):

maybe,

OpenStudy (anonymous):

At first I thought it wasn't c or d but I have a lot of stuff going on in my head so I could change my opinion at any moment now. Lol

OpenStudy (sleepyjess):

Lol I know it isn't A or C because if you put 100 into the equation it doesn't equal less than 100.

OpenStudy (anonymous):

Ohhh

OpenStudy (anonymous):

Ok but now that i think deeply about it, I think that it is D because i don't think a decimal would work well with that equation..

OpenStudy (sleepyjess):

But it would still work. \(\large52+2(23.8)=52+47.6=99.6\)

OpenStudy (anonymous):

so all real numbers?

OpenStudy (sleepyjess):

\(\huge\frak\color{#57ff14}{Correct}\)

OpenStudy (anonymous):

thank you!